How many moles of HCl are present in 40 mL of a .035 M solution?
1
Expert's answer
2020-09-01T07:36:08-0400

Number of moles"=" Molarity "\times" Volume ( in Litres)

Volume "( V) = 40\ ml = \frac{40}{1000} \ L= 0.04 \ L"

Molarity "( M) = 0.35\ M"

Number of moles "=0.04\times 0.35=0.014"